# Querying Autonomous Vehicle Point Clouds: Enhanced by 3D Object Counting with CounterNet

### v1

@@ Abstract.
Point cloud, retrieval, query, autonomous vehicle
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Computer vision problems
@@ 1. Introduction
The rise of autonomous driving has significantly increased the importance of point cloud data analysis, as autonomous vehicles heavily rely on point clouds to perceive their surroundings accurately. While vast amounts of data are collected—3D detection alone can generate approximately 5TB of point cloud data per hour (Kazhamiaka et al., 2021)—not all data is relevant to every training task. For instance, detecting vehicle collisions and analyzing traffic congestion differ in their focus. Collision detection often involves identifying sudden, localized changes in vehicle count within short time frames, while congestion analysis requires aggregating vehicle counts over longer durations to identify sustained traffic buildup. These tasks, therefore, depend on different object types, time scales, and aggregation strategies, reinforcing the need for flexible point cloud analytics designed for specific use cases.
@@
These query types form the foundation of our approach. However, their effectiveness relies heavily on accurate object counts within individual frames. Unfortunately, current 3D detection models struggle to provide reliable counts across object categories, as we show in Figure 2. This undermines the assumption made by many existing query systems that detection outputs can serve as ground truth.
Prior work mostly focused on 2D traffic video querying(Kang et al., 2022, 2019; Bang et al., 2023; Cao et al., 2022). These studies generally assume that 2D object detection provides ground truth semantic information, and focus primarily on building efficient and effective indexing mechanisms to reduce the cost of query execution. For example, BlazeIt (Kang et al., 2019) and TASTI (Kang et al., 2022) use proxy models to filter simple scenes, Seiden (Bang et al., 2023) uses sampling to minimize model invocations, and Figo (Cao et al., 2022) dynamically selects detectors based on video complexity. Although fewer works address 3D point cloud querying, approaches like MAST (Li et al., 2025) extend Seiden (Bang et al., 2023) to the 3D domain by incorporating the spatial-temporal filters into the indexing process. These methods inherit a key assumption from their 2D counterparts: that object detection models yield ground truth results to serve as a foundation for query execution.
@@
We propose CounterNet, a heatmap-based baseline designed for accurate object counting in complex traffic scenes. Our approach focuses on enhancing semantic information extraction from point cloud frames, with a particular emphasis on achieving precise, category-aware object counts. To improve its adaptability across object types and sizes, we refine CounterNet to support multi-class counting through scale-aware mechanisms. In addition, we introduce a per-frame dynamic model selection strategy, which assigns the most suitable counting model configuration to each frame based on its complexity, further boosting overall query accuracy without sacrificing efficiency. It is worth noting that our study’s primary focus is the query accuracy rather than efficiency.
@@
  • To support query accuracy, we introduce CounterNet, a heatmap-based network that serves as our baseline solution. CounterNet focuses on detecting object centers, enabling accurate and efficient object counting within large-scale traffic point cloud frames (Section 4).
@@ 2. Related Work
Video Data Querying and Indexing. Most of the related studies have focused on video data querying. Benefiting from advances in object detection in images, such as ResNet (He et al., 2016), Mask R-CNN (He et al., 2017), and YOLO (Jocher et al., 2020), recent studies on video querying has primarily focused on indexing techniques to enhance query efficiency. For example, studies (Kang et al., 2022, 2019; Cao et al., 2022) utilise proxy models for fast indexing. To be specific, Kang et al. (2019) introduced a specialized neural network as a similarity model, employing a compact ResNet architecture. Kang et al. (2022) proposed a low-cost embedding deep neural network (DNN) that was trained using triplet loss. Cao et al. (2022) focused on the ensemble of various models applied to different video segments, with the primary task being the strategic segmentation of video frames and matching segments with appropriate query models. In the most recent work, SEIDEN (Bang et al., 2023) constructed the index by effectively sampling the most important frames, allowing only these selected frames to be processed by the detection model for semantic information extraction, considering the high cost of running detection models. Additionally, SEIDEN leveraged object information extracted by the detection model to linearly predict the contents of non-sampled frames, enabling full indexing.
@@
Point Cloud Data Querying and Indexing. Building on SEIDEN’s framework, Li et al. (2025) extends this approach to traffic point cloud data. It further enhances sampling by incorporating reinforcement learning with a spatially informed reward function and predicts the contents of non-sampled frames using spatial-temporal information extracted from sampled frames via the detection model. Similar to video query solutions, these studies primarily focus on improving query efficiency through optimized indexing. However, they often overlook the detection model’s actual performance in extracting semantic information from point cloud frames.
3D Object Detection in Point Clouds. 3D object detection employs a backbone for feature extraction and a detection head for task-specific outputs. Early works such as PointNet (Qi et al., 2017a), PointNet++ (Qi et al., 2017b), and VoxelNet (Zhou and Tuzel, 2018) laid the foundation for 3D feature extraction. Later advancements (Liu et al., 2021; Zhou et al., 2020b; Lai et al., 2023) introduced improved feature extraction methods. Detection heads in 3D systems follow two primary approaches: anchor-based (Yan et al., 2018; Lang et al., 2019; Shi et al., 2020; Yang et al., 2020; Chen et al., 2017) and center-based (Zhou et al., 2020a; Chen et al., 2023; Bai et al., 2022; Sun et al., 2022).
Despite substantial progress, these detection models are not designed for querying tasks. Detection focuses on object identification, whereas querying emphasizes extracting task-specific information, requiring tailored solutions.
Object Counting in Images and Point Cloud. The object count within video frames is crucial for supporting queries. Numerous studies have addressed object counting (Chattopadhyay et al., 2017; Stahl et al., 2018; Cholakkal et al., 2020; Laradji et al., 2018; Xu et al., 2023), noting that counting becomes significantly harder as the number of objects increases. Recent research has utilized textual information to improve accuracy (Jiang et al., 2023; Kang et al., 2024). However, 3D data presents greater challenges than 2D due to the irregular distribution of points in point cloud frames, unlike the fixed grids of 2D images.
While some studies have investigated 3D object counting (Jenkins et al., 2023; Lesani et al., 2020; Nellithimaru and Kantor, 2019), they often focus on simpler scenarios like pedestrian or crop counting. These methods struggle with the complexity of autonomous vehicle scenes, which feature diverse object types in a single frame.
@@ 3. Problem Definition
Autonomous vehicles generate extensive point cloud data, which is crucial for navigation and environmental understanding. Since onboard systems prioritize real-time processing, they usually have limited storage. Instead, point cloud data is transmitted to a remote, e.g., edge server, enabling centralized processing and analysis(Liu et al., 2019; Tang et al., 2021; Ming, 2023). Following Liu et al. (Liu et al., 2019), the data collected by the vehicle is uploaded to the cloud, and a NoSQL database is used to store the extracted features for querying. To achieve this, we first define the query types.
@@ 4. Method: CounterNet
Data Model. Given that object count is the most important information for querying (Kang et al., 2022, 2019; Cao et al., 2022; Bang et al., 2023), to accommodate the distributed nature of point cloud data collected by autonomous vehicles, instead of directly following the SQL-like relational database (Dönderler et al., 2005; Le et al., 2008; Lu et al., 2015), we define a NoSQL data model as the document-like structure. The detailed definition can be found in Appendix C.1.
@@
Design Motivation of CounterNet. The most straightforward method for obtaining object counts is through a regression model. However, a single point cloud frame contains vast amounts of information, including multiple objects in varying numbers, making it difficult for a regression model to capture such fine-grained details. Alternatively, object detection models are commonly used for counting, but they focus heavily on localizing objects, such as drawing accurate bounding boxes and determining object orientation. This complexity causes them to underperform in tasks focused solely on object counting.
@@
It motivated us to develop a new method, which has the following properties: (1) To mitigate the limitation in the regression model, the method should be capable of detecting the appearance of the object in the region of the point cloud for a more fine-grained detection. (2) To mitigate the limitation brought by the detection model, the method should concentrate on detecting the appearance of object instead of spending efforts on the localization of objects.
@@
Network Architecture. We propose to utilise heatmaps for object counting, as illustrated in Figure 3. The objective of this approach is to detect object centers, which are represented as peaks in the heatmap. This method satisfied the two expected property mentioned above: (1) Compared to simple regression, the heatmap emphasizes identifying object appearances by detecting centers, resulting in more accurate counts. (2) It simplifies the detection task by disregarding object size and orientation, focusing exclusively on detecting the object centers to identify their presence.
@@
Heatmap. With the projected feature map, we generate a K-channel heatmap ($K=|C|$), where each channel corresponds to a specific object class. To train the heatmap, we utilise the ground truth positions of objects, annotating object centers with higher values and surrounding areas with lower values, which we utilise the Gaussian Kernal for a smooth (Zhou et al., 2020a; Law and Deng, 2018; Yin et al., 2021).
@@
Object Counting via Threshold-based Peak Detection. We count objects by finding local peaks in the heatmap. Those peaks are detected object centers. Specifically, we apply a 2D local maxima algorithm  [^1] https://mathworks.com/help/matlab/ref/islocalmax2.html to identify object centers, which then allows us to estimate both the count and approximate positions of the objects. Thresholding is crucial for reducing noise. To suppress noise in low-activation regions of heatmaps, we apply a fixed threshold to the predicted heatmap before applying local maxima detection. This ensures that only confident peaks, corresponding to likely object centers, are retained. The threshold is defined during the training stage. Since the heatmap values are normalized within the range of 0 to 1, we select a mid-range value as the threshold to ensure effective filtering (i.e., 0.5). Intuitively, the model is expected to assign higher values (above 0.5) to regions corresponding to object grids, while non-object regions will have lower values (below 0.5).
@@
We use thresholded local maxima on the heatmap to derive the object count. To supervise this process, we introduce an L1 loss for counting accuracy as shown in Equation 2. In this equation, $y$ represents the predicted count, while $\hat{y}$ represents the ground truth count. The count loss provides an additional supervisory signal to the heatmap, ensuring that regions corresponding to object centers exceed the threshold while other regions remain below it. Additionally, while the focal loss provides direct supervision for the heatmap, the count loss serves as auxiliary supervision for the entire network, guiding it toward more precise object counting results.
@@ 5. Performance Optimizations of CounterNet
To further enhance CounterNet’s performance, we introduce a partition-based approach that divides the point cloud into smaller regions, improving counting accuracy (in Section 5.1). Additionally, we propose adding overlap between partitions to more accurately detect objects near the partition boundaries, leading to more precise and reliable counting results (in Section 5.2). We also provide a model selection method to dynamically select the model trained with different parameters to further optimize the counting accuracy across different object type (in Section 5.3).
@@ 5.1. CounterNet with Feature Map Partition
To address this, we apply Otsu’s method (Otsu et al., 1975) during the inference stage to dynamically determine the optimal threshold $k$ for each partition.
Otsu’s method maximizes the between-class variance, given by the formula: $\alpha_{B}^{2}(k)=\omega_{0}(k)\times\omega_{1}(k)\times\left[mean_{0}(k)-mean_{1}(k)\right]^{2}$. Here, $\omega_{0}(k)$ and $\omega_{1}(k)$ are the proportion of points below and above the threshold, respectively, while $mean_{0}(k)$ and $mean_{1}(k)$ represent the mean values of the points below and above $k$, respectively. By maximizing this function, we find the optimal threshold $k$ for each partition. The computed threshold $k$ is then used to distinguish relevant and irrelevant regions within the partition. We combine this dynamic threshold $k$ with the fixed threshold $t$ in training. If $k\geq t$, we adopt $k$ as the threshold for partition, otherwise, we use $t$ as the threshold.
Loss. Since partitioning the feature map introduces imbalance across partitions. To mitigate this, we apply weights to the count loss for each partition, as shown in Equation 3, allowing the model to focus more on high-density partitions.
@@ 5.2. Feature Map Partition with Overlaps
An intuitive approach to prevent objects from being split across partitions is to use feature-based clustering methods such as K-Means (Faber et al., 1994). However, clustering presents several challenges in this context. First, it is computationally expensive and struggles to differentiate objects in high-density frames, where object boundaries may be unclear. Second, clustering does not provide guarantees on the shape of partitions, which can result in irregular regions that disrupt backpropagation during training. These limitations make clustering an impractical solution for integrating partitioning into deep learning models.
To overcome these challenges, we propose a simpler and more effective alternative by expanding partition regions to introduce overlap (we use $o$ to represent the expanding ratio), as illustrated in Figure 5. This approach reduces the likelihood of splitting an object across partitions by ensuring that if an object’s center falls near a partition boundary, it is also captured within the central region of an adjacent partition. Compared to clustering, this method has several advantages: (1) it is lightweight and adds minimal computational overhead while preserving the benefits of partitioning, and (2) it integrates seamlessly into the CounterNet pipeline, maintaining consistency across frames.
The training and inference details for CounterNet with overlapped partition can be found in Appendix D.1.

### v2

@@ Abstract.
Point Cloud, Aggregation Query, Autonomous Vehicle
copyright: acmlicensedjournalyear: 2025doi: 10.1145/3746027.3755043conference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Ireland.booktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelandisbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Computer vision problems
@@ 1. Introduction
The rise of autonomous driving has significantly increased the importance of point cloud data analysis, as autonomous vehicles heavily rely on point clouds to perceive their surroundings accurately. While vast amounts of data are collected, 3D detection alone can generate approximately 5TB of point cloud data per hour (Kazhamiaka et al., 2021), not all data is relevant to every training task. For instance, detecting vehicle collisions and analyzing traffic congestion differ in their focus. Collision detection often involves identifying sudden, localized changes in vehicle count within short time frames, while congestion analysis requires aggregating vehicle counts over longer durations to identify sustained traffic buildup. These tasks, therefore, depend on different object types, time scales, and aggregation strategies, reinforcing the need for flexible point cloud analytics designed for specific use cases.
@@
These query types form the foundation of our approach. However, their effectiveness relies heavily on accurate object counts within individual frames. Unfortunately, current 3D detection models struggle to provide reliable counts across object categories, as shown in Figure 2. This undermines the assumption made by many existing query systems that detection outputs can serve as ground truth.
Prior work mostly focused on 2D traffic video querying(Kang et al., 2022, 2019; Bang et al., 2023; Cao et al., 2022). These studies generally assume that 2D object detection provides ground truth semantic information, and focus primarily on building efficient and effective indexing mechanisms to reduce the cost of query execution. For example, BlazeIt (Kang et al., 2019) and TASTI (Kang et al., 2022) used proxy models to filter simple scenes, Seiden (Bang et al., 2023) used sampling to minimize model invocations, and Figo (Cao et al., 2022) dynamically selected detectors based on video complexity. Although fewer works address 3D point cloud querying, approaches like MAST (Li et al., 2025) extended Seiden (Bang et al., 2023) to the 3D domain by incorporating the spatial-temporal filters into the indexing process. These methods made a key assumption from their 2D counterparts: that object detection models yield ground truth results to serve as a foundation for query execution.
@@
We propose CounterNet, a heatmap-based baseline designed for accurate object counting in complex traffic scenes. Our approach focuses on enhancing semantic information extraction from point cloud frames, with a particular emphasis on achieving precise, category-aware object counts. To improve its adaptability across object types and sizes, we refine CounterNet to support multi-class counting through scale-aware mechanisms. In addition, we introduce a per-frame dynamic model selection strategy, which assigns the most suitable counting model configuration to each frame based on its complexity, further boosting overall query accuracy without sacrificing efficiency. It is worth noting that our study’s primary focus is query accuracy, rather than efficiency.
@@
  • To support query accuracy, we introduce CounterNet, a heatmap-based network that serves as our baseline solution. CounterNet focuses on detecting object centers, enabling object counting within large-scale traffic point cloud frames (Section 4).
@@ 2. Related Work
We review the literature from the most relevant domains: video and point cloud data querying and indexing. We further extend it to less related areas, such as 3D object detection and object counting.
Video Data Querying and Indexing. Most of the related studies have focused on video data querying. Benefiting from advances in object detection in images, such as ResNet (He et al., 2016), Mask R-CNN (He et al., 2017), and YOLO (Jocher et al., 2020), recent studies on video querying have primarily focused on indexing techniques to enhance query efficiency. For example, studies (Kang et al., 2022, 2019; Cao et al., 2022) utilised proxy models as the index for efficient querying. To be specific, Kang et al. (2019) introduced a specialized neural network as a similarity model, employing a compact ResNet architecture. Kang et al. (2022) proposed a low-cost embedding deep neural network (DNN) that was trained using triplet loss. Cao et al. (2022) focused on the ensemble of various models applied to different video segments, with the primary task being the strategic segmentation of video frames and matching segments with appropriate query models. In the most recent work, SEIDEN (Bang et al., 2023) constructed the index by effectively sampling the most important frames, allowing only these selected frames to be processed by the detection model for semantic information extraction, considering the high cost of running detection models. Additionally, SEIDEN leveraged object information extracted by the detection model to linearly predict the contents of non-sampled frames, enabling indexing for all frames.
@@
Point Cloud Data Querying and Indexing. Building on SEIDEN’s framework, Li et al. (2025) extended this approach to traffic point cloud data. It further enhanced sampling by incorporating reinforcement learning with a spatially informed reward function and predicted the contents of non-sampled frames using spatial-temporal information extracted from sampled frames via the detection model. Similar to video query solutions, these studies primarily focused on improving query efficiency through optimized indexing. However, they often overlooked the detection model’s actual performance in extracting semantic information from point cloud frames.
3D Object Detection in Point Clouds. 3D object detection employs a backbone for feature extraction and a detection head for task-specific outputs. Early works such as PointNet (Qi et al., 2017a), PointNet++ (Qi et al., 2017b), and VoxelNet (Zhou and Tuzel, 2018) laid the foundation for 3D feature extraction. Later advancements (Liu et al., 2021; Zhou et al., 2020b; Lai et al., 2023) introduced improved feature extraction methods. Detection heads in 3D systems follow two primary approaches: anchor-based (Yan et al., 2018; Lang et al., 2019; Shi et al., 2020; Yang et al., 2020; Chen et al., 2017) and center-based (Zhou et al., 2020a; Chen et al., 2023; Bai et al., 2022; Sun et al., 2022). Despite substantial progress, these detection models were not designed for querying tasks. Detection focused on object identification, whereas querying emphasized extracting task-specific information.
Object Counting in Images and Point Cloud. The object counts within video frames is crucial for supporting queries. Numerous studies have addressed object counting (Chattopadhyay et al., 2017; Stahl et al., 2018; Cholakkal et al., 2020; Laradji et al., 2018; Xu et al., 2023), noting that counting becomes significantly harder as the number of objects increases. Recent research utilized textual information to improve accuracy (Jiang et al., 2023; Kang et al., 2024). However, 3D data presented greater challenges than 2D due to the irregular distribution of points in point cloud frames, unlike the fixed grids of 2D images. While some studies investigated 3D object counting (Jenkins et al., 2023; Lesani et al., 2020; Nellithimaru and Kantor, 2019), they often focused on simpler scenarios like pedestrian or crop counting. These methods struggled with the complexity of autonomous vehicle scenes, which feature diverse object types in a single frame.
@@ 3. Problem Definition
Autonomous vehicles generate extensive point cloud data, which is crucial for environmental understanding. Since onboard systems prioritize real-time processing, they usually have limited storage. Instead, point cloud data is transmitted to a remote, e.g., edge server, enabling centralized processing and analysis(Liu et al., 2019; Tang et al., 2021; Ming, 2023). Following Liu et al. (Liu et al., 2019), the data collected by the vehicle is uploaded to the cloud, and a NoSQL database is used to store the extracted features for querying. To achieve this, we first define the query types.
@@ 4. Method: CounterNet
Data Model. Given that object count is the most important information for querying (Kang et al., 2022, 2019; Cao et al., 2022; Bang et al., 2023), we define a NoSQL data model as the document-like structure. The detailed definition can be found in Appendix C.1.
@@
Design Motivation of CounterNet. The most straightforward method for obtaining object counts is through a regression model. However, a single point cloud frame contains vast amounts of information, including multiple objects in varying numbers, making it difficult for a regression model to capture such fine-grained details. Alternatively, object detection models are focused on localizing objects, such as drawing accurate bounding boxes and determining object orientation. This complexity causes them to underperform in tasks focused solely on object counting.
@@
This motivates us to develop a method that has the following properties: (1) To mitigate the limitation in the regression model, the method should be capable of detecting the appearance of the object in the region of the point cloud for a more fine-grained detection. (2) To mitigate the limitation brought by the detection model, the method should concentrate on detecting the appearance of objects instead of spending efforts on the localization of objects.
@@
Network Architecture. We propose to utilise heatmaps for object counting, as illustrated in Figure 3. The objective of this approach is to detect object centers, which are represented as peaks in the heatmap. This method satisfies the two expected property mentioned above: (1) Compared to simple regression, the heatmap emphasizes identifying object appearances by detecting centers, resulting in more accurate counts. (2) It simplifies the detection task by disregarding object size and orientation, focusing exclusively on detecting the object centers to identify their presence.
@@
Heatmap. With the projected feature map, we generate a K-channel heatmap ($K=|C|$), where each channel corresponds to a specific object type. To train the heatmap, we utilise the ground truth positions of objects, annotating object centers with higher values and surrounding areas with lower values (Zhou et al., 2020a; Law and Deng, 2018; Yin et al., 2021).
@@
Object Counting via Threshold-based Peak Detection. We count objects by finding local peaks in the heatmap. These peaks correspond to detected object centers. Specifically, we apply a 2D local maxima algorithm  [^1] https://mathworks.com/help/matlab/ref/islocalmax2.html to identify object centers, which then allows us to estimate both the count and approximate positions of the objects. Thresholding is crucial for reducing noise. To suppress noise in low-activation regions of heatmaps, we apply a fixed threshold to the predicted heatmap before applying local maxima detection. This ensures that only confident peaks, corresponding to likely object centers, are retained. The threshold is determined during the training stage. Since the heatmap values are normalized within the range of 0 to 1, we select a mid-range value as the threshold to ensure effective filtering (i.e., 0.5). Intuitively, the model is expected to assign higher values to regions corresponding to object grids, while non-object regions will have lower values.
@@
We use thresholded local maxima on the heatmap to derive the object count. To supervise this process, we introduce an L1 loss for counting accuracy as shown in Equation 2. In this equation, $y$ represents the predicted count, while $\hat{y}$ represents the ground truth count. The count loss provides an additional supervisory signal to the heatmap, ensuring that regions corresponding to object centers exceed the threshold while other regions remain below it.
@@ 5. Performance Optimizations of CounterNet
To further enhance CounterNet’s performance, we introduce a partition-based approach that divides the point cloud into smaller regions, improving counting accuracy (in Section 5.1). Additionally, we propose adding overlap between partitions to more accurately detect objects near the partition boundaries, leading to more precise and reliable counting results (in Section 5.2). We also provide a model selection method to dynamically select the model trained with different parameters to further optimize the counting accuracy across different object types (in Section 5.3).
@@ 5.1. CounterNet with Feature Map Partition
To address this, we apply Otsu’s method (Otsu et al., 1975) during the inference stage to dynamically determine the optimal threshold $k$ for each partition. Otsu’s method maximizes the between-class variance, given by the formula: $\alpha_{B}^{2}(k)=\omega_{0}(k)\times\omega_{1}(k)\times\left[mean_{0}(k)-mean_{1}(k)\right]^{2}$. Here, $\omega_{0}(k)$ and $\omega_{1}(k)$ are the proportion of points below and above the threshold, respectively, while $mean_{0}(k)$ and $mean_{1}(k)$ represent the mean values of the points below and above $k$, respectively. By maximizing this function, we find the optimal threshold $k$ for each partition. The computed threshold $k$ is then used to distinguish relevant and irrelevant regions within the partition. We combine this dynamic threshold $k$ with the fixed threshold $t$ in training. If $k\geq t$, we adopt $k$ as the threshold for partition; otherwise, we use $t$ as the threshold.
Loss. Partitioning the feature map introduces imbalance across partitions. To mitigate this, we apply weights to the count loss for each partition, as shown in Equation 3, allowing the model to focus more on high-density partitions.
@@ 5.2. Feature Map Partition with Overlaps
An intuitive approach to prevent objects from being split across partitions is to use feature-based clustering methods such as K-Means (Faber et al., 1994). However, clustering presents several challenges in this context. First, it is computationally expensive and struggles to differentiate objects where object boundaries may be unclear. Second, clustering produces irregular regions that disrupt backpropagation during training. These limitations make clustering an impractical solution for integrating with our pipelines.
To overcome these challenges, we propose a simpler and more effective alternative by expanding partition regions to introduce overlap, as illustrated in Figure 5. This approach reduces the likelihood of splitting an object across partitions by ensuring that if an object’s center falls near a partition boundary, it is also captured within the central region of an adjacent partition. Compared to clustering, this method has several advantages: (1) it is lightweight and adds minimal computational overhead while preserving the benefits of partitioning, and (2) it integrates seamlessly into the CounterNet pipeline, maintaining consistency across frames.
The training and inference details is in Appendix D.1.